# Universal Definitions of the Roman Factorial: Introduction to Foundational Functions and the Generalization Process

### v2

@@ 1.1 Contents
The contents of this study are divided into 5 parts, each containing several sections, which are further subdivided into subsections. For example, you are currently reading Part 1, Section 1, Subsection 1.1.
Part 1 contains four sections. Sections 3 and 5 are about foundational functions (F.F.) while Sections 4 and 6 are about redefining the Roman factorial recursively and directly [^1] The terms ”recursively” and ”directly,” as well as the concept of ”foundational functions,” will be clarified shortly. Any unfamiliar mathematical terminology introduced in this section or later will be explained in time.. This paper focuses on introducing the F.F. step by step, in a way that builds intuition and insight.
@@
Later parts of this study include applications of this generalization process to the double factorial, triple factorial and all higher order factorials, as well as complex factorials and factorials of non-integer order. The next part specifically is about multifactorial extensions of the Roman factorial.
Additionally, for those interested in diving deeper into specific mathematical concepts, an addendum is provided in Section 9. Lastly, Section 8 lists all references used in the formulation of this paper along with other details and information.
@@ 1.2 Useful information
Next, you’ll find a list of key words and phrases essential for understanding this paper:
@@
  • Roman factorial: An expansion of the factorial to negative integers.
@@
  • Direct relationship: A mathematical definition that directly expresses a sequence or function without referring to its own previous values. Also: non-recursive, iterative
@@
  • F.F.: Foundational functions, a series of simple Boolean-like functions that are built upon each other and have mostly binary outputs (0 or 1).
@@ 2.1 Factorial
$3!=3\cdot 2\cdot 1=6$
$4!=4\cdot 3\cdot 2\cdot 1=24$
$5!=5\cdot 4\cdot 3\cdot 2\cdot 1=120$
@@
$n!=\prod_{k=1}^{n}k\enspace,\quad n\in\mathbb{N}\thinspace.$ (Eq. 2.1.1)
@@
$n!=\prod_{k=0}^{n-1}(n-k)\enspace,\quad n\in\mathbb{N}\thinspace.$ (Eq. 2.1.2)
@@
We see that $0!=1$. This is a convention for the empty product [^4] More info about the empty product in Addendum 9.2. and also because it extends various combinatoric identities. For more reasons why 0 factorial should equal 1 see Addendum 9.7, but we will show a few of them shortly.
@@
$3!=3\cdot 2!=3\cdot 2\cdot 1$
$4!=4\cdot 3!=4\cdot 3\cdot 2\cdot 1$
$5!=5\cdot 4!=5\cdot 4\cdot 3\cdot 2\cdot 1$
@@
$n!=n(n-1)!$
We can reiterate this property to redefine the factorial recursively, like so:
$n!=n(n-1)!\thinspace,\quad 0!=1\thinspace,\quad n\in\mathbb{N}\thinspace.$ (Eq. 2.1.3)
Eq. 2.1.3 is called the recursive definition of the factorial, because in order to calculate $n!$ we need to know $(n-1)!\thinspace$, starting with $0!=1$. Using this relationship we can find the next larger factorial for a given factorial. For example, if we know what $8!$ is, we can easily find $9!$ by evaluating for $n=9$: $9!=9\cdot 8!=362880$.
@@
If we substitute $n$ with $n+1$ in Eq. 2.1.3 and solve for $n!\thinspace$, we get a similar equation [^5] In this paper, the symbol $\mathbb{N}$ refers to natural numbers or positive integers $(1,2,3,\dots)$ while $\mathbb{N}_{0}$ means non-negative integers $(0,1,2,3,\dots)$. More about number sets in Addendum 9.1.:
$n!=\cfrac{(n+1)!}{n+1}\enspace,\quad n\in\mathbb{N}_{0}\thinspace.$ (Eq. 2.1.4)
@@
$3!=\cfrac{4\cdot 3\cdot 2\cdot 1}{4}=\cfrac{4!}{4}$
$2!=\cfrac{3\cdot 2\cdot 1}{3}=\cfrac{3!}{3}$
$1!=\cfrac{2\cdot 1}{2}=\cfrac{2!}{2}$
$1\displaystyle 0!=\cfrac{1}{1}=$ 1
@@
When trying to find factorials of negative numbers, for example $(-1)!\thinspace$, we come across a problem. We initially notice that there is no intuitive definition of the factorial for negative integers.
Let’s approach this by attempting to find $(-1)!$ recursively. Using the recurrence relationship in Eq. 2.1.4 for $n=-1$, we observe that we have to divide by 0:
$(-1)!=\cfrac{(-1+1)!}{-1+1}=\cfrac{0!}{0}=\cfrac{1}{0}\enspace,$
@@
$0!=0\cdot(-1)!\enspace\Rightarrow\enspace 1=0\cdot(-1)!\thinspace,$
which is impossible since there is no value of $(-1)!$ that satisfies this equation.
@@ 2.2 Roman factorial
The Roman factorial [1] [2] is an extension of the factorial function to negative integers. It is named after Steven Roman, who used it in his work about umbral calculus. Roman factorial can be expressed in the following closed form:
$\lfloor n\rceil!=\begin{cases}\thinspace n!&,\enspace n\in\mathbb{N}_{0}\\ \cfrac{(-1)^{-n-1}}{(-n-1)!}&,\enspace n\in\mathbb{Z}^{-},\end{cases}$ (Eq. 2.2.1)
@@
$n!=n(n-1)!\thinspace,\quad 0!=1\thinspace,\quad n\in\mathbb{N}\thinspace.$ (Eq. 2.2.2)
The variable $n$ can take only integer values, as denoted by $n\in\mathbb{Z}$. The following table is a list of negative integer factorials, as defined by Eq. 2.2.1:
@@
Negative Roman factorials exhibit a pattern akin to the reciprocals of positive factorials, with alternating signs. Notably, there is an offset: $(-5)!$ is not the reciprocal of $5!$, but rather of $4!$. However, this offset does not affect any properties of the factorial.
@@
$*.\lfloor n\rceil!=\cfrac{\lfloor n+1\rceil!}{n+1}\enspace,\enspace n\in\mathbb{Z}^{*}.$ (Eq. 2.2.3)
This relationship is defined for any integer $n$ except 0, which is expressed mathematically by the set $\mathbb{Z}$ with an asterisk. Eq. 2.2.3 does not define the Roman factorial recursively, it is just an identity valid for all integers.
@@
$\lfloor n\rceil!=\begin{cases}n\lfloor n-1\rceil!&,\enspace n\in\mathbb{N}\\[3.68027pt] 1&,\enspace n=\{0\thinspace,-1\}\\[1.18027pt] \cfrac{\lfloor n+1\rceil!}{n+1}&,\enspace n\in\mathbb{Z}^{-}\setminus\{-1\}\thinspace,\end{cases}$ (Eq. 2.2.4)
where $\mathbb{Z}^{-}\setminus\{-1\}$ stands for the set [^6] More about number sets on Addendum 9.1. of all negative integers except $-1$.
@@
$\lfloor-2\rceil!=\cfrac{\lfloor-2+1\rceil!}{-2+1}=\cfrac{\lfloor-1\rceil!}{-1}=-1\thinspace.$
@@
Certainly, there remains an anomaly in Eq. 2.2.3 particularly at $n=0$, yet the piece-wise Roman factorial offers a solution: $\rceil!=1$ 1. This specific value doesn’t resolve the division by 0; indeed, no number can. Instead, the recursive relationship fails to provide a value for $(-1)!$ or 1$\lfloor$-1$\rceil$!⌊ - 1 ⌉ !, so we address this limitation by asserting that the relationship is undefined at this anomaly (assigning $(-1)!$ the value of 1 arbitrarily).
Let’s examine an example of generalizing Eq. 2.2.4 so that it becomes a single expression, universally defined across its domain. To begin with, consider the following Boolean-valued functions [^7] More about Boolean functions in Addendum 9.5.:
$f_{1}\equiv f_{1}(n)=\begin{cases}+1\thinspace,&n\geq 0\\ -1\thinspace,&n<0\thinspace,\end{cases}$ (Eq. 2.2.5)
$f_{2}\equiv f_{2}(n)=\begin{cases}0\thinspace,&n\geq 0\\ +1\thinspace,&n<0\thinspace.\end{cases}$ (Eq. 2.2.6)
@@
$\begin{gathered}\lfloor n\rceil!=\big{(}n+f_{2}\big{)}^{f_{1}}\lfloor n-f_{1}\rceil!\thinspace,\enspace n\neq\{0,-1\}\thinspace,\\[6.18027pt] \lfloor 0\rceil!=\lfloor-1\rceil!=1\thinspace,\enspace n\in\mathbb{Z}\thinspace.\end{gathered}$ (Eq. 2.2.7)
This result will be explained thoroughly in Section 4 and serves as an example of the generalization process used in this paper. This procedure seeks to rewrite various closed-form definitions as singular expressions by employing a set of foundational functions which have simple and analytic forms.
@@
The non-recursive definition of the Roman factorial will have the form
$\lfloor n\rceil!=F(n)\thinspace,\quad n\in\mathbb{Z}\thinspace,$ (Eq. 2.2.8)
where $F(n)$ stands for a product-like function, similar to Eq. 2.1.2 or Eq. 2.1.1 for positive integers.
The recursive expression for the Roman factorial will be based on Eq. 2.2.4 or Eq. 2.2.1 but it will be a singular expression, universally defined across all integers.
@@ 3.1 Introduction
These so-called foundational functions (F.F.) are simple and are built upon each other, beginning with one called $\delta(n)$. They take integer or real values of $n$ and are categorized based on their outputs, which are similar to those of Boolean-valued functions. Information about Boolean algebra is provided in Addendum 9.5.
@@
Now, let’s introduce the concept of output tables. Whenever a new function is defined in this paper, there will be a table showcasing its output values. For instance, the function $\sin(x)$ is represented by the following output table:
$x$ 0 $\nicefrac{{\pi}}{{2}}$ $\pi$ $\nicefrac{{3\pi}}{{2}}$ $2\pi$
$\sin(x)$ 0 1 0 -1 0
Table Tbl. 3.1.1: The function $\sin(x)$
@@
The values of $n$ that will be used extensively are 0 and any positive or negative numbers. It should be noted that all F.F. will converge at the limits as $n$ goes to infinity, unlike $\sin(x)$. Most of the time, foundational functions will also have the same value across all positive or negative numbers.
@@
This function can be expressed by the output pattern $[0,\nicefrac{{1}}{{e}},1,\nicefrac{{1}}{{e}},0]$. Most commonly though, output tables will have less columns. We can condense Tbl. 3.1.2 like so:
@@
In the rest of Section 3, as well as in Section 5, we will define a set of functions denoted by (F.F.). They will be helpful in Section 4 where we will express Eq. 2.2.1 as a single relationship and in Section 6 in which we will build two universal product definitions for the Roman factorial. Foundational functions will be necessary for intuitively comprehending the mathematical definitions presented in this paper.
@@ 3.2 The function $\delta(n)$
The first function to be defined here is denoted $\delta(n)$. For any integer $n$, $\delta(n)$ takes the form:
$\delta(n)=n+\varepsilon\thinspace,\quad 0<\varepsilon<1\thinspace,\quad n\in\mathbb{Z}\thinspace.$ (Eq. 3.2.1)
@@
$\delta(n)=\begin{cases}+\thinspace,\quad n\geq 0\\ -\thinspace,\quad n<0\end{cases},\enspace n\in\mathbb{Z}\thinspace.$ (Eq. 3.2.2)
@@
$\delta(n)=\lfloor n\rfloor+\varepsilon\thinspace,\quad 0<\varepsilon<1\thinspace,\quad n\in\mathbb{R}\thinspace.$ (Eq. 3.2.3)
@@
$\lfloor 1.3\rfloor=1\qquad\lfloor 2.7\rfloor=2$
$\lfloor 3\rfloor=3\qquad\lfloor-3.4\rfloor=-4$
@@
Eq. 3.2.3 is termed the universal definition of $\delta(n)$. Unlike Eq. 3.2.2, this formulation is not piece-wise and accommodates inputs from all real numbers [^8] For a more advanced definition of $\delta(n)$ involving a Fourier approximation, see Addendum 9.4..
@@
As observed, discontinuities arise naturally due to the floor function. Indeed, there is no continuous definition of $\delta(n)$ because of Bolzano’s Theorem: for a continuous $\delta(n)$, it would be $\delta(n)=0$ for some value between $\delta(-1)=-1+\varepsilon<0$ and $\delta(1)=\varepsilon>0$, contradicting the requirement that $\delta(n)$ should not output zero for any input.
This function serves as a fundamental component for all subsequent F.F. presented in this paper. A well-defined $\delta(n)$ for all real numbers is imperative, even if non-integers are not utilized in this paper.

### v7

# References.bib

@@ 1 Preface
Universal Definitions of the Roman Factorial is a study that expands the concept of the factorial to various domains and develops mathematical expressions to describe it.
The study is too large and extensive to be contained in a single paper, so it is divided in parts. Part 1 is considered an introduction to its follow-ups, but it stands alone as a analysis into the Roman factorial.
@@ 1.1 Contents
The contents of this study are divided into 5 parts, of which this paper is the first one. Each part contains several sections, and they are further subdivided into subsections. For example, you are currently reading Part 1, Section 1, Subsection 1.1.
Part 1 contains four main sections. Sections 3 and 5 are about foundational functions (F.F.) while Sections 4 and 6 are about redefining the Roman factorial recursively and non-recursively [^1] The term ”recursively”, as well as the concept of ”foundational functions”, will be clarified shortly. Any unfamiliar mathematical terminology introduced in this section or later will be explained in time.. This paper focuses on introducing the F.F. step by step, in a way that builds intuition and insight.
@@
Later parts of this study [^2] The next part of this study will introduce a notation system for the different types of factorial variations. However, since this paper focuses exclusively on the Roman factorial, there’s no need to distinguish between similar concepts here. Therefore, what might later be called $n!_{(1)}$ or $n!_{1}$ will simply be referred to as $\lfloor n\rceil!$ in this paper. include applications of this generalization process to the double factorial, triple factorial and all higher order factorials. There is a part about generalizing the factorial to non-integers, and another about non-integer orders. The last is about complex factorials and a new formulation of the F.F. presented in this paper to accommodate for the new domain.
Concepts like those mentioned in the paragraph above will not be explained and analyzed here, so there is no need to learn about them now. In fact, no prior knowledge is required about any advanced mathematical concept, as this paper was written with accessibility in mind. Anyone with high-school education should be able to read this study all the way through, which is its main goal in the first place.
For those interested in diving deeper into specific mathematical concepts, an addendum is provided in Section 9. It has basic information about a variety of things and is intended to fill any gaps not explicitly elaborated upon in the main text. Lastly, Section 8 lists all references used in the formulation of this paper along with other details and information.
@@ 1.2 Useful information
Equations or tables may be repeated throughout this paper, but their original numbering is preserved. For instance, Eq. 2.2.1 appears many times in different sections, but each time it has the same tag. It is also clickable as a hyperlink, referencing its first appearance in the text.
This paper uses terminology that may be unfamiliar at first. For this reason, a list of key words and phrases are provided here:
@@
  • Roman factorial: An extension of the factorial into negative integers.
@@
  • Non-recursive relationship: A mathematical definition that directly expresses a sequence or function without referring to its own previous values. Also: direct, iterative
@@
  • F.F.: Foundational functions, a series of simple Boolean-like functions that are built upon each other and usually have binary outputs (0 or 1).
@@ 2.1 Factorial
$3!=3\cdot 2\cdot 1=6\,,$
$4!=4\cdot 3\cdot 2\cdot 1=24\,,$
$5!=5\cdot 4\cdot 3\cdot 2\cdot 1=120\,,$
@@
$n!=\prod_{k=1}^{n}k\,,\quad n\in\mathbb{Z}^{+}.$ (Eq. 2.1.1)
@@
$n!=\prod_{k=0}^{n-1}(n-k)\,,\quad n\in\mathbb{Z}^{+}.$ (Eq. 2.1.2)
@@
We see that $0!=1$. This is a convention for the empty product [^5] More info about the empty product in Addendum 9.2. and also because it extends various combinatoric identities. For more reasons why $0!$ should equal 1 see Addendum 9.7, but we will show a few of them shortly.
@@
$3!=3\cdot 2!=3\cdot 2\cdot 1\,,$
$4!=4\cdot 3!=4\cdot 3\cdot 2\cdot 1\,,$
$5!=5\cdot 4!=5\cdot 4\cdot 3\cdot 2\cdot 1\,,$
@@
$n!=n(n-1)!\,.$
We can include the starting value $0!=1$ to define the factorial recursively, like so:
$n!=n(n-1)!\,,\quad 0!=1\,,\quad n\in\mathbb{Z}^{+}.$ (Eq. 2.1.3)
Eq. 2.1.3 is called the recursive definition of the factorial, because in order to calculate $n!$ we need to know $(n-1)!\,$, starting with $0!=1$. Using this relationship we can find the next larger factorial for a given factorial. For example, if we know what $8!$ is, we can easily find $9!$ by evaluating for $n=9$: $9!=9\cdot 8!=362880$.
@@
If we substitute $n$ with $n+1$ in Eq. 2.1.3 and solve for $n!\,$, we get a similar equation [^6] In this paper, the symbol $\mathbb{Z}^{+}$ refers to natural numbers or positive integers $(1,2,3,\dots)$ while $\mathbb{Z}^{+}_{0}$ means non-negative integers $(0,1,2,3,\dots)$. More information about number sets in Addendum 9.1.:
$n!=\cfrac{(n+1)!}{n+1}\>,\quad n\in\mathbb{Z}^{+}_{0}\,.$ (Eq. 2.1.4)
@@
$3!=\cfrac{4\cdot 3\cdot 2\cdot 1}{4}=\cfrac{4!}{4}\>,$
$2!=\cfrac{3\cdot 2\cdot 1}{3}=\cfrac{3!}{3}\>,$
$1!=\cfrac{2\cdot 1}{2}=\cfrac{2!}{2}\>,$
$0!=\cfrac{1}{1}=1\>.$
@@
When trying to find factorials of negative numbers, for example $(-1)!\,$, we come across a problem. We initially notice that there is no intuitive definition of the factorial for negative integers, so we need a new definition entirely or at least an expansion of the recursive relationship as iterated in Eq. 2.1.4.
Let’s approach this issue by attempting to find $(-1)!$ recursively. Using the recurrence relationship in Eq. 2.1.4 for $n=-1$, we observe that we have to divide by 0:
$(-1)!=\cfrac{(-1+1)!}{-1+1}=\cfrac{0!}{0}=\cfrac{1}{0}\>,$
@@
$0!=0\cdot(-1)!\enspace\Rightarrow\enspace 1=0\cdot(-1)!\,,$
which is impossible since there is no value of $(-1)!$ that satisfies this equation (no number multiplied by 0 gives 1).
@@ 2.2 Roman factorial
The Roman factorial [roman] [romanfact] is an extension of the factorial function to negative integers [^7] The symbols [roman] and [romanfact] are references to citations, listed in Section 8.. It is named after Steven Roman, who used it in his work about umbral calculus. Roman factorial can be expressed in the following closed form:
$\lfloor n\rceil!=\begin{cases}\,n!&,\enspace n\in\mathbb{Z}^{+}_{0}\\ \cfrac{(-1)^{-n-1}}{(-n-1)!}&,\enspace n\in\mathbb{Z}^{-},\end{cases}$ (Eq. 2.2.1)
@@
$n!=n(n-1)!\,,\quad 0!=1\,,\quad n\in\mathbb{Z}^{+}.$ (Eq. 2.1.3)
The variable $n$ can take only integer values, as denoted by the sets $\mathbb{Z}^{+}_{0}$ and $\mathbb{Z}^{-}$. The following table is a list of negative integer Roman factorials, as defined by Eq. 2.2.1:
@@
Negative Roman factorials exhibit a pattern akin to the reciprocals of positive factorials, with alternating signs. Notably, there is an offset: 5$\lfloor$-5$\rceil$!⌊ - 5 ⌉ ! is not the reciprocal of $5!$, but rather of $4!$. However, this offset does not affect any properties of the factorial.
@@
$\left\lfloor n\right\rceil!=\cfrac{\left\lfloor n+1\right\rceil!}{n+1}\,,\quad n\in\mathbb{Z}^{-}\setminus\{-1\}.$ (Eq. 2.2.2)
This relationship is defined for any integer $n$ except -1, which is expressed mathematically by the set [^8] More about number sets on Addendum 9.1. $\mathbb{Z}^{-}\setminus\{-1\}$. Eq. 2.2.2 does not define the Roman factorial recursively, it is just an identity valid for all integers except $-1$.
@@
$\lfloor n\rceil!=\begin{cases}n\lfloor n-1\rceil!&,\enspace n\in\mathbb{Z}^{+}\\[3.68027pt] 1&,\enspace n=\{0\,,-1\}\\[1.18027pt] \cfrac{\lfloor n+1\rceil!}{n+1}&,\enspace n\in\mathbb{Z}^{-}\setminus\{-1\}\,.\end{cases}$ (Eq. 2.2.3)
@@
$\lfloor-2\rceil!=\cfrac{\lfloor-2+1\rceil!}{-2+1}=\cfrac{\lfloor-1\rceil!}{-1}=-1\,.$
@@
Certainly, there remains an anomaly in Eq. 2.2.2 particularly at $n=-1$, yet the piece-wise Roman factorial offers a solution: $\rceil!=1$ 1. This specific value doesn’t resolve the division by 0; indeed, no number can. Instead, the recursive relationship fails to provide a value for $(-1)!$ or 1$\lfloor$-1$\rceil$!⌊ - 1 ⌉ !, so we address this limitation by asserting that the relationship is undefined at this anomaly (assigning $(-1)!$ the value of 1 arbitrarily).
In these last few paragraphs of the introduction, we will show an example of generalizing Eq. 2.2.3 so that it becomes a single expression, universally defined across its domain. To begin with, consider the following Boolean-valued functions [^9] More about Boolean functions in Addendum 9.5.:
$f_{1}\equiv f_{1}(n)=\begin{cases}+1\,,&n\geq 0\\ -1\,,&n<0\,,\end{cases}$ (Eq. 2.2.4)
$f_{2}\equiv f_{2}(n)=\begin{cases}0\,,&n\geq 0\\ +1\,,&n<0\,.\end{cases}$ (Eq. 2.2.5)
@@
$\begin{gathered}\lfloor n\rceil!=\big{(}n+f_{2}\big{)}^{f_{1}}\lfloor n-f_{1}\rceil!\,,\enspace n\neq\{0,-1\}\,,\\[6.18027pt] \lfloor 0\rceil!=\lfloor-1\rceil!=1\,,\quad n\in\mathbb{Z}\,.\end{gathered}$ (Eq. 2.2.6)
The process of reaching this result will be explained thoroughly in Section 4 and it serves as an example of the generalization process used in this paper. This procedure seeks to rewrite various closed-form definitions as singular expressions by employing a set of foundational functions which have simple and analytic forms.
@@
Lastly, the non-recursive definition of the Roman factorial will have the form
$\lfloor n\rceil!=F(n)\,,\quad n\in\mathbb{Z}\,,$ (Eq. 2.2.7)
where $F(n)$ stands for a product-like function, similar to Eq. 2.1.1 or Eq. 2.1.2 for positive integers. The recursive expression for the Roman factorial will be based on Eq. 2.2.3 or Eq. 2.2.1 but it will be a singular expression, universally defined across all integers.
@@ 3.1 Introduction
These so-called foundational functions (F.F.) are simple and are built upon each other, beginning with one called $\delta(n)$. They take integer or real values of $n$ and are categorized based on their outputs, which are similar to those of Boolean-valued functions. Basic information about Boolean algebra is provided in Addendum 9.5.
@@
Now, let’s introduce the concept of output tables. Whenever a new function is defined in this paper, there will be a table showcasing its output values. For instance, the function $\sin\theta$ is represented by the following output table:
$\theta$ 0 $\nicefrac{{\pi}}{{2}}$ $\pi$ $\nicefrac{{3\pi}}{{2}}$ $2\pi$
$\sin\theta$ 0 1 0 -1 0
Table Tbl. 3.1.1: The function $\sin\theta$
@@
The values of $n$ that will be used extensively are 0 and any positive or negative numbers. It should be noted that all F.F. will converge at the limits as $n$ goes to infinity, unlike $\sin\theta$. Most of the time, foundational functions will also have the same value across all positive or negative numbers.
@@
This function can be expressed by the output pattern $[0,\nicefrac{{1}}{{e}},1,\nicefrac{{1}}{{e}},0]$. Most commonly though, output tables will have less columns. Using the previous example, we can condense Tbl. 3.1.2 like so:
@@
In the rest of Section 3, as well as in Section 5, we will define a set of functions denoted by (F.F.). They will be helpful in Section 4 where we will condense Eq. 2.2.1 and Eq. 2.2.3 into a single relationship each, and in Section 6 in which we will build two universal product definitions for the Roman factorial. Foundational functions will be necessary for intuitively comprehending the mathematical definitions presented in this paper.
@@ 3.2 The function $\delta(n)$
The first function to be defined here is named $\delta(n)$. For any integer $n$, $\delta(n)$ takes the form:
$\delta(n)=n+\varepsilon\,,\quad 0<\varepsilon<1\,,\quad n\in\mathbb{Z}\,.$ (Eq. 3.2.1)
@@
$\delta(n)=\begin{cases}+\,,\quad n\geq 0\\ -\,,\quad n<0\end{cases},\enspace n\in\mathbb{Z}\,.$ (Eq. 3.2.2)
@@
$\delta(n)=\lfloor n\rfloor+\varepsilon\,,\quad 0<\varepsilon<1\,,\quad n\in\mathbb{R}\,.$ (Eq. 3.2.3)
@@
$\lfloor 1.3\rfloor=1\,,\qquad\lfloor 2.7\rfloor=2\,,$
$\lfloor 3\rfloor=3\,,\qquad\lfloor-3.4\rfloor=-4\,.$
@@
Eq. 3.2.3 is termed the universal definition of $\delta(n)$. Unlike Eq. 3.2.2, this formulation is not piece-wise and accommodates inputs from all real numbers.
@@
As observed, discontinuities arise naturally due to the floor function. Indeed, there is no continuous definition with $\delta(n)$’s outputs because of Bolzano’s Theorem: for a continuous $\delta(n)$, it would be $\delta(n)=0$ for some value between $\delta(-1)=-1+\varepsilon<0$ and $\delta(1)=\varepsilon>0$, contradicting the requirement that $\delta(n)$ should not output zero for any input.
This function serves as a fundamental component for all subsequent F.F. presented in this paper. A well-defined $\delta(n)$ for all real numbers is imperative [^10] For a more advanced definition of $\delta(n)$ involving a Fourier approximation, see Addendum 9.4., even if non-integers are not utilized in this paper.